prove that sin (degree(1)) is an algebraic number


1
Expert's answer
2022-03-18T01:09:08-0400

180o= "\pi" radian

Let 1 be k

sin ko

ko= k"\pi"/180

ek"\pi"i/180 = cos ( k"\pi"/180)+ isin (k"\pi" /180)

(ek"\pi"i/180)180 = cos (k"\pi") + i sin (k"\pi")

(-1)k =+1= -1

cos (k"\pi"/180) + i sin (k"\pi"/180)

cos (k"\pi"/180) - i sin (k"\pi"/180)


2cos (k"\pi"/180) is an algebraic integer

cos (k"\pi"/180) becomes an algebraic integer

i sin (k"\pi"/180) is an algebraic integer

sin (k"\pi"/180) is also an algebraic integer


"\therefore" sin (degree(1)) is an algebraic number